OpenStudy (anonymous):

From a shipment of 300 light-bulbs, a sample of 60 was selected at random and tested. If 18 light-bulbs in the sample were found to be defective, how many defective light-bulbs would be expected in the entire shipment?

OpenStudy (anonymous):

90 I think

OpenStudy (anonymous):

18/60=.3 or 30% So P(Defective item from shipment)=.30 (Equation 1) This is the probability equals 30 percent But P(Defective item from shipment)=n/300 (Equation 2) where n is the number of defective light bulbs over the entire population, here it is equal to 300 Now set Equation 2 (right hand side) equal to Equation 1 (right hand side) n/300=.30 n=.30*300=90 (multiplied left and right hand side of the above equation by 300 So that means you expect to have 90 defective light bulbs in the entire shipment

OpenStudy (anonymous):

Really there are many different ways to explain this but it depends what class it is for (there are many probability classes out there)...

OpenStudy (goformit100):

90 IS CORRECT